OpenStudy (anonymous):

Which of the following is the solution to the equation 8^(y – 2) = 2^y ?

OpenStudy (anonymous):

y = 3, y = -1, y = –3, and y = 1

OpenStudy (anonymous):

2^(3(y - 2)) = 2^y 2^(3y - 6) = 2^y 3y - 6 = y 2y = 6 y = 3

OpenStudy (anonymous):

Option 1.

OpenStudy (lalaly):

\[8^{y-2} = \frac{8^y}{8^2}\]\[\frac{1}{8^2} (2^3)^y = 2^y\]

OpenStudy (lalaly):

pratus way is easier xD